An investment promises to pay $30 each year for the next four years. In four years, it will also pay $1000. Your required return on this investment is 7%. What is the value of the investment today? What is the value of the investment in one year? What is the value of the investment in two years? What is the value of the investment in ten years?
Answer #1

The present value of the investment is calcuated as follows

- $30_- $30_- $30_1 S1030 . =(1+0.07)*(1+0.07) ? *(1+0.07)3*(1+0.07) 4

The value of the investment today = $ 864.51

Future value of investment = Present value \times ( 1 + interest rate)n

Value of the investment in one year = $ 864.51 \times (1 + 0.07)1  = $ 925.03

Value of the investment in two years = $ 864.51 \times (1 + 0.07)2   = $ 989.78

Value of the investment in ten years =   $ 864.51 \times (1 + 0.07)10   = $ 1700.62